# FlowLog: Efficient and Extensible Datalog via Incrementality

### v1

@@ 1. Introduction
The rapid expansion of data-intensive applications has underscored the need for query languages that are both simple and expressive. As a declarative language, Datalog adds recursion to relational algebra in a concise syntax, making it especially well-suited for domains such as graph processing (Lam et al., 2013; Fan et al., 2019), network monitoring (Abiteboul et al., 2005), program analysis (Scholz et al., 2016; Szabó et al., 2016), and distributed systems (Chu et al., 2024). A classic illustration is the graph reachability query, which determines all nodes reachable from a source node a. Its Datalog formulation is:
@@
In recent years, academic advances and industry demands have spurred the development of Datalog engines. In program analysis, Datalog has been proven powerful for static analysis tasks such as bug detection and security checks. This momentum has led to systems such as Soufflé (Scholz et al., 2016), Flix (Madsen et al., 2016), Flan (Abeysinghe et al., 2024) and Ascent (Sahebolamri et al., 2022), which are designed from the ground up, often augmented by domain-specific optimizations (Subotic et al., 2018; Jordan et al., 2019). Although these systems excel in their intended applications, they tend to sacrifice flexibility. For example, adding incremental maintenance for continuous updates—highly desirable for prototyping and debugging of Datalog applications—typically requires major system modifications (Zhao et al., 2021). Similarly, many of these systems were initially designed in single-node architectures; scaling up/out to accommodate larger datasets frequently demands extensive engineering.
@@
Other systems sidestep the from-scratch approach by layering Datalog functionality atop established databases for out-of-the-box deployment. A prominent example is RecStep (Fan et al., 2019), which compiles Datalog into SQL queries, then delegates its execution to QuickStep (Patel et al., 2018)—a single-node in-memory parallel dbms—one iteration at a time. RecStep inherits mature database features from QuickStep, such as the cost-based optimizer and parallelism. However, reusing a black-box dbms complicates Datalog-specific optimizations, particularly those spanning multiple iterations, such as semi-naïve evaluation and index maintenance. While RecStep interjects between iterations to impose some of these optimizations, the back-and-forth control flow incurs non-negligible synchronization overheads. A similar design underpins DDlog (Ryzhyk and Budiu, 2019), an incremental-only Datalog engine that directly translates Datalog into Differential Dataflow [^1] Differential dataflow (McSherry et al., 2013) programs chain up a set of DD’s streaming operators that continuously maintain internal states for efficient incremental computation as data evolves (see Sec. 2.3 for a formal introduction.) (DD) programs. Despite its simplicity, studies have observed that DDlog incurs significant memory overhead–often orders of magnitude higher than alternative approaches (Mansur et al., 2023; Zhao et al., 2021; Hu et al., 2021; Li et al., 2022).
@@
Balancing flexibility and efficiency for Datalog remains an open challenge. This paper addresses this by pursuing a new design that unifies $(i)$ efficient, off-the-shelf relational operators as execution primitives, and $(ii)$ flexible, fine-grained optimization controls. To achieve the first goal, we reuse DD’s streaming operators as building blocks, positioning DDlog as a baseline that directly code-generates Datalog into lower-level DD without a distinct optimization phase. Achieving the second goal requires a novel design for Datalog optimization. There are two main reasons for this: $(i)$ the primary bottleneck of DDlog is memory usage, and $(ii)$ recursion weakens many conventional SQL techniques—e.g., cost-based planning (Leis et al., 2015a) lacks reliable static statistics in recursive contexts—so existing systems (e.g., Soufflé, DDlog) fall back on ad hoc heuristics or manual performance tuning. To curb memory usage, we apply a suite of memory-focused rewrites on a relational intermediate representation (IR) that fully separates the rule’s logical plan from its physical realization. Additionally, instead of the conventional cost-based planning, we opt for robustness-aware optimizations, which prioritizes avoiding worst-case scenarios (e.g., pathological join orders and large intermediate relations) over seeking a best query plan.
@@
Related Work. As (Fan et al., 2019; Ketsman and Koutris, 2022) pointed out, most Datalog engines are purpose-built for specific domains (Scholz et al., 2016; Hu et al., 2021; Szabó et al., 2016; Madsen et al., 2016; Sahebolamri et al., 2022; Abeysinghe et al., 2024; Herlihy et al., 2024; Shkapsky et al., 2016; Sun et al., 2025). Over time, they have introduced a range of optimizations and extensions tailored to Datalog (Hu et al., 2021; Green et al., 2012; Shaikhha et al., 2024; Wu et al., 2022; Khamis et al., 2023; Zhang et al., 2023; Gilray et al., 2024), such as incremental Datalog (Zhao et al., 2021; Ryzhyk and Budiu, 2019). We incorporate some of these techniques into FlowLog (e.g., index sharing (Subotic et al., 2018; McSherry et al., 2020), unified idb evaluation (Fan et al., 2019)), while others, such as magic sets (Wang et al., 2022), de-specialized relations (Hu et al., 2021), customized data structures (Sahebolamri et al., 2023) and worst-case optimal joins (Abeysinghe et al., 2024), remain promising future explorations. Notably, major gaps in the Datalog literature persist, particularly the lack of effective query planning and limited scalability in highly iterative workloads, both being critical challenges for large-scale Datalog applications (Arch et al., 2022; Fan et al., 2022).
@@ 2.1. Datalog Basics
A standard Datalog program (Abiteboul et al., 1995) is a set of rules. A rule is an expression of the following form:
@@
Common Datalog Extensions. To enrich Datalog for practical usage, we incorporate common syntactic extensions as (Fan et al., 2019)–constraints, stratified negations (where negated atoms are either edb or an idb from a lower stratum), and (possibly recursive) aggregations. For example, this allows finding two hops $(x,z)$ in a graph that are:
@@ 2.3. Differential Dataflow
Differential Dataflow (or DD(McSherry et al., 2013) is a data-parallel programming model for large-scale incremental data processing. Its Rust implementation [^2] https://github.com/TimelyDataflow/differential-dataflow compiles down to Timely Dataflow, a lower-level generic distributed streaming system introduced by  (Murray et al., 2013).
@@
Arrangements. An arrangement is an in-memory index for DD collections (McSherry et al., 2020). It can be considered as a sorted dictionary that allows efficient concurrent access. An arrangement indexes batches of historical changes, maintains them over time, and merges them into compact representations as appropriate (e.g., when a timestamp is advanced). The first join in the above code pre-arranges both operands by imposing an arrange operation internally for each and then uses a more primitive join_core operator to join on arrangements. Indeed, the first join is executed under the hood as
@@ 3. System Design
In this section, we present the architecture of FlowLog. Like DDlog, FlowLog uses DD as its execution backend. However, rather than compiling Datalog directly into a DD program, FlowLog first translates the rules into a more amenable intermediate representation (IR) that represents the logical structure of the input program, before lowering into DD. In this design, we can reason about optimizations only at the logical level by abstracting away many low-level complexities underneath DD, such as recursive control, de-duplications, and incrementality, making it simpler to modularize and extend. The IR is akin to the recent idea of grounding, which was used to unravel efficient algorithms for Datalog (Khamis et al., 2023; Zhao et al., 2024). We incorporate some of these theoretical ideas when optimizing the IR (see Section 5).
@@ 4. Logic Fusion
Multiple Concat. A Concat transformation unions two input collections as one. In Datalog, there are often multiple rules deriving the same idb. Instead of unioning two at a time and maintaining every intermediate Concat, we fuse multiple into a single but multi-way ConcatAll transformation to union all inputs at once. It compiles down to a physical concat_all operator. This is analogous to the unified idb evaluation optimization proposed in RecStep (Fan et al., 2019).
@@ 5. Join-project Plan Optimization
Traditional dbms optimizers rely on approximate statistics to choose plans at the lowest cost (Leis et al., 2015b). However, finding optimal query plans for Datalog rules is much more strenuous as these statistics are often missing or unstable: idb (s) accumulates at runtime in varying delta sizes across iterations, and real-world applications such as DOOP (Bravenboer and Smaragdakis, 2009) and DDISASM (Flores-Montoya and Schulte, 2020) often involve highly complex join topologies (e.g. cyclic joins) than those typically handled by traditional dbms. Hence systems Soufflé and DDlog give up and use only hard-coded listing orders: if a rule is written with body $R,S,T$, the system executes joins in that order even if there are cross products, i.e. $(R\bowtie S)\bowtie T$. RecStep collects runtime statistics periodically and invokes the dbms’s optimizer to re-optimize query plans on the fly, paying a synchronization and catalog maintenance overhead.
FlowLog’s optimizer chooses a static join-project plan for every Datalog rule; and maps it one-to-one to the IR. Rather than relying on any runtime statistics, it analyzes the join graph of a rule via a hybrid of conventional heuristics (e.g. filter pushdown) and worst-case-aware analysis (Zhao et al., 2024). While this approach is simple and oriented toward worst-case guarantees, our experiments (Sec. 10) demonstrate that it often avoids disastrous join orders, even if it does not always identify the optimal one. This provides a principled foundation for developing fine-grained cost models for Datalog, for example, by incorporating edb statistics in the future.
@@ 5.1. Structural Cost Model
We define the cost of a join-project plan (or equivalently, an IR) as the maximum cost of any transformation it contains. In Figure 2(b), the maximum is 3, determined by the lower Join-FlatMap. Intuitively, for Example 2.1 over an input graph of $n$ nodes, this plan implies a worst-case intermediate size (and so is the asymptotic runtime) of $O(n^{3})$. The same reasoning extends to arbitrary Datalog programs: the structural cost upper-bounds worst-case intermediate output sizes, and formal guarantees are in (Zhao et al., 2024). Similar techniques are proposed for subgraph pattern matching (McMahan et al., 2004) and have shown effectiveness for multi-way many-to-many joins.
@@ 5.2. Search Strategy
A join tree (Yannakakis, 1981) is a spanning tree of the join graph such that for every variable $x$, the atoms containing $x$ (in its schema) induce a connected subtree of the spanning tree. The join tree is then used to define acyclicity of a rule, i.e. a rule is acyclic if and only if there is a join tree for it. Zhao et al. (Zhao et al., 2024) showed that for acyclic rules, bottom-up join orders (rooted arbitrarily) along a join tree, after early projections, usually yield tight intermediate size bounds.
A join spanning tree (JST) extends this notion to cyclic rules. A JST is a maximum spanning tree of the weighted join graph and reduces to a join tree when the rule is acyclic (Maier, 1983). A rooted JST defines a join-project plan by post-order traversal: at each step, an atom is joined with its parent, followed by projecting out variables no longer needed. The only JST (also a join graph or a join tree) for Example 2.1 is shown in Figure 3. Here we root the JST at $\textsf{edge}(x,y)$, and it maps into the IR on the right, i.e. a bottom-up join order. If we root the JST at $\textsf{reach}(z)$ instead, we recover the IR in Figure 2(b).
@@
Search Space. Our search space excludes semijoins (atoms whose variables are subsumed by another), antijoins and filters, as they are pushed down to the lowest possible transformation in the IR. This narrows down to a multi-way (inner) join where the search space is defined as all rooted JSTs of its join graph. There are three reasons for such a choice: (1) JSTs avoid cross products when possible, since cross products correspond to zero-weight edges in the weighted join graph, (2) this space is reasonably small and easy to enumerate (Winter, 1986), and (3) it collapses down to rooted join trees for acyclic rules, for which the post-order join-project plans yield tight bounds on the intermediate sizes (proven by (Zhao et al., 2024)). In our running example (which is acyclic), Figure 3 will be selected against Figure 2(b) as it has a cost of 2, instead of 3 in our cost model. Intuitively, it avoids computing $\textsf{edge}(x,y)\bowtie\textsf{edge}(y,z)$ by using two semi-joins.
@@
Consider an expensive rule (Figure 4, left) from the DOOP program analysis framework (Bravenboer and Smaragdakis, 2009), which contains a recursive 8-way join. VarType, HeapAllocationType, and ComponentType are edbs; others (in bold) are idbs. Figure 4(center) shows its join graph. The JST selected by the optimizer is rooted at LoadArrayIdx and is annotated on the join graph by thick, directed edges. The dotted edges are edges in the join graph, but are not part of the rooted JST. Here, Reach(inm) is a semijoin atom to be subsumed by LoadArrayIdx and inm is projected away after semijoin. When creating the IR, this semijoin is pushed down to the leaf LoadArrayIdx.
@@ 5.3. Plan Execution
Left-deep Plans. Join planning is tied closely to the underlying execution. The listing order Soufflé, DDlog and others used is equivalent to a left-deep join plan in a dbms and is incapable of representing bushy plans, such as the one in Figure 4. This is a necessary choice for Soufflé as it always compiles the listing order into an indexed nested loop join, where indexes (e.g. B-trees) are built on every atom except the first one and it iterates over each tuple of the first atom (as the outer loop) and probes into the indexes of the rest. Left-deep plans allow pipelining without intermediate materialization, which makes Soufflé memory-efficient. However, scaling indexed nested loop joins to a multi-threaded setting is non-trivial (e.g. (Leis et al., 2014)) and modern Datalog systems such as Soufflé and Flan (Abeysinghe et al., 2024) only distribute the outermost for-loop among threads. As shown in our experiments, this level of parallelism is insufficient to saturate resources even for simple recursive queries like reachability.
Bushy Plans. FlowLog targets DD, a fundamentally different execution model. Its differential operators are inherently stateful, so intermediate materialization is unavoidable. A key advantage, however, is their asynchrony: changes propagate through the dataflow graph and operators react concurrently without waiting for earlier stages to finish. Hence, DD naturally exploits multicore parallelism, as computation is scheduled dynamically based on available updates rather than a pre-determined execution sequence. Consequently, FlowLog can take advantage of bushy plans (i.e. IR is bushy) without worrying about compromising parallelism. To control memory blow-ups, FlowLog leverages insights from database theory (e.g. tree decompositions (Khamis et al., 2016, 2017; Zhao et al., 2024), worst-case optimal joins (Ngo et al., 2013; Wang et al., 2023), etc.) to select plans that have smallest possible worst-case intermediate sizes. When multiple candidate plans have the same estimated cost, FlowLog heuristically prefers bushier (shallower) rooted JSTs.

### v3

@@ 1. Introduction
The rapid expansion of data-intensive applications has underscored the need for query languages that are both simple and expressive. As a declarative language, Datalog adds recursion to relational algebra in a concise syntax, making it especially well-suited for domains such as graph processing (SociaLite13; recstep), network monitoring (Abiteboul05), program analysis (Bernhard2016; Szabo2016), and distributed systems (Chu2024). A classic illustration is the graph reachability query, which determines all nodes reachable from a source node a. Its Datalog formulation is:
@@
In recent years, academic advances and industry demands have spurred the development of Datalog engines. In program analysis, Datalog has been proven powerful for static analysis tasks such as bug detection and security checks. This momentum has led to systems such as Soufflé (Bernhard2016), Flix (Flix16), Flan (Flan24) and Ascent (Ascent), which are designed from the ground up, often augmented by domain-specific optimizations (IndexSouffle; TrieSouffle). Although these systems excel in their intended applications, they tend to sacrifice flexibility. For example, adding incremental maintenance for continuous updates—highly desirable for prototyping and debugging of Datalog applications—typically requires major system modifications (David2021). Similarly, many of these systems were initially designed in single-node architectures; scaling up/out to accommodate larger datasets frequently demands extensive engineering.
@@
Other systems sidestep the from-scratch approach by layering Datalog functionality atop established databases for out-of-the-box deployment. A prominent example is RecStep (recstep), which compiles Datalog into SQL queries, then delegates its execution to QuickStep (QuickStep)—a single-node in-memory parallel dbms—one iteration at a time. RecStep inherits mature database features from QuickStep, such as the cost-based optimizer and parallelism. However, reusing a black-box dbms complicates Datalog-specific optimizations, particularly those spanning multiple iterations, such as semi-naïve evaluation and index maintenance. While RecStep interjects between iterations to impose some of these optimizations, the back-and-forth control flow incurs non-negligible synchronization overheads. A similar design underpins DDlog (ddlog), an incremental-only Datalog engine that directly translates Datalog into Differential Dataflow [^1] Differential dataflow (McSherryMII13) programs chain up a set of DD’s streaming operators that continuously maintain internal states for efficient incremental computation as data evolves (see Sec. 2.3 for a formal introduction.) (DD) programs. Despite its simplicity, studies have observed that DDlog incurs significant memory overhead–often orders of magnitude higher than alternative approaches (MansurWC23; David2021; souffleInterp; dyck22).
@@
Balancing flexibility and efficiency for Datalog remains an open challenge. This paper addresses this by pursuing a new design that unifies $(i)$ efficient, off-the-shelf relational operators as execution primitives, and $(ii)$ flexible, fine-grained optimization controls. To achieve the first goal, we reuse DD’s streaming operators as building blocks, positioning DDlog as a baseline that directly code-generates Datalog into lower-level DD without a distinct optimization phase. Achieving the second goal requires a novel design for Datalog optimization. There are two main reasons for this: $(i)$ the primary bottleneck of DDlog is memory usage, and $(ii)$ recursion weakens many conventional SQL techniques—e.g., cost-based planning (LeisGMBK015) lacks reliable static statistics in recursive contexts—so existing systems (e.g., Soufflé, DDlog) fall back on ad hoc heuristics or manual performance tuning. To curb memory usage, we apply a suite of memory-focused rewrites on a relational intermediate representation (IR) that fully separates the rule’s logical plan from its physical realization. Additionally, instead of the conventional cost-based planning, we opt for robustness-aware optimizations, which prioritizes avoiding worst-case scenarios (e.g., pathological join orders and large intermediate relations) over seeking a best query plan.
@@
Related Work. As (recstep; DBLP:journals/ftdb/KetsmanK22) pointed out, most Datalog engines are purpose-built for specific domains (Bernhard2016; souffleInterp; Szabo2016; Flix16; Ascent; Flan24; DBLP:conf/icde/HerlihyMAO24; DBLP:conf/sigmod/ShkapskyYICCZ16; GPUDatalog; shovon2025multi). Over time, they have introduced a range of optimizations and extensions tailored to Datalog (souffleInterp; LogicBlox; NestedRec24; DCDatalog22; datalogo; egglog; 10.14778/3712221.3712232), such as incremental Datalog (David2021; ddlog). We incorporate some of these techniques into FlowLog (e.g., index sharing (IndexSouffle; arrangement), unified idb evaluation (recstep)), while others, such as magic sets (WangK0PS22), de-specialized relations (souffleInterp), customized data structures (DBLP:journals/pacmpl/SahebolamriBMM23) and worst-case optimal joins (Flan24), remain promising future explorations. Notably, major gaps in the Datalog literature persist, particularly the lack of effective query planning and limited scalability in highly iterative workloads, both being critical challenges for large-scale Datalog applications (DBLP:conf/lopstr/ArchHZSS22; DBLP:conf/datalog/FanMK22).
@@ 2.1. Datalog Basics
A standard Datalog program (abiteboul1995foundations) is a set of rules. A rule is an expression of the following form:
@@
Common Datalog Extensions. To enrich Datalog for practical usage, we incorporate common syntactic extensions as (recstep)–constraints, stratified negations (where negated atoms are either edb or an idb from a lower stratum), and (possibly recursive) aggregations. For example, this allows finding two hops $(x,z)$ in a graph that are:
@@ 2.3. Differential Dataflow
Differential Dataflow (or DD(McSherryMII13) is a data-parallel programming model for large-scale incremental data processing. Its Rust implementation [^2] https://github.com/TimelyDataflow/differential-dataflow compiles down to Timely Dataflow, a lower-level generic distributed streaming system introduced by  (Timely).
@@
Arrangements. An arrangement is an in-memory index for DD collections (arrangement). It can be considered as a sorted dictionary that allows efficient concurrent access. An arrangement indexes batches of historical changes, maintains them over time, and merges them into compact representations as appropriate (e.g., when a timestamp is advanced). The first join in the above code pre-arranges both operands by imposing an arrange operation internally for each and then uses a more primitive join_core operator to join on arrangements. Indeed, the first join is executed under the hood as
@@ 3. System Design
In this section, we present the architecture of FlowLog. Like DDlog, FlowLog uses DD as its execution backend. However, rather than compiling Datalog directly into a DD program, FlowLog first translates the rules into a more amenable intermediate representation (IR) that represents the logical structure of the input program, before lowering into DD. In this design, we can reason about optimizations only at the logical level by abstracting away many low-level complexities underneath DD, such as recursive control, de-duplications, and incrementality, making it simpler to modularize and extend. The IR is akin to the recent idea of grounding, which was used to unravel efficient algorithms for Datalog (datalogo; grounding24). We incorporate some of these theoretical ideas when optimizing the IR (see Section 5).
@@ 4. Logic Fusion
Multiple Concat. A Concat transformation unions two input collections as one. In Datalog, there are often multiple rules deriving the same idb. Instead of unioning two at a time and maintaining every intermediate Concat, we fuse multiple into a single but multi-way ConcatAll transformation to union all inputs at once. It compiles down to a physical concat_all operator. This is analogous to the unified idb evaluation optimization proposed in RecStep (recstep).
@@ 5. Join-project Plan Optimization
Traditional dbms optimizers rely on approximate statistics to choose plans at the lowest cost (DBLP:journals/pvldb/LeisGMBK015). However, finding optimal query plans for Datalog rules is much more strenuous as these statistics are often missing or unstable: idb (s) accumulates at runtime in varying delta sizes across iterations, and real-world applications such as DOOP (BravenboerS09) and DDISASM (DDISASM) often involve highly complex join topologies (e.g. cyclic joins) than those typically handled by traditional dbms. Hence systems Soufflé and DDlog give up and use only hard-coded listing orders: if a rule is written with body $R,S,T$, the system executes joins in that order even if there are cross products, i.e. $(R\bowtie S)\bowtie T$. RecStep collects runtime statistics periodically and invokes the dbms’s optimizer to re-optimize query plans on the fly, paying a synchronization and catalog maintenance overhead.
FlowLog’s optimizer chooses a static join-project plan for every Datalog rule; and maps it one-to-one to the IR. Rather than relying on any runtime statistics, it analyzes the join graph of a rule via a hybrid of conventional heuristics (e.g. filter pushdown) and worst-case-aware analysis (grounding24). While this approach is simple and oriented toward worst-case guarantees, our experiments (Sec. 10) demonstrate that it often avoids disastrous join orders, even if it does not always identify the optimal one. This provides a principled foundation for developing fine-grained cost models for Datalog, for example, by incorporating edb statistics in the future.
@@ 5.1. Structural Cost Model
We define the cost of a join-project plan (or equivalently, an IR) as the maximum cost of any transformation it contains. In Figure 2(b), the maximum is 3, determined by the lower Join-FlatMap. Intuitively, for Example 2.1 over an input graph of $n$ nodes, this plan implies a worst-case intermediate size (and so is the asymptotic runtime) of $O(n^{3})$. The same reasoning extends to arbitrary Datalog programs: the structural cost upper-bounds worst-case intermediate output sizes, and formal guarantees are in (grounding24). Similar techniques are proposed for subgraph pattern matching (joinwidth) and have shown effectiveness for multi-way many-to-many joins.
@@ 5.2. Search Strategy
A join tree (Yannakakis81) is a spanning tree of the join graph such that for every variable $x$, the atoms containing $x$ (in its schema) induce a connected subtree of the spanning tree. The join tree is then used to define acyclicity of a rule, i.e. a rule is acyclic if and only if there is a join tree for it. Zhao et al. (grounding24) showed that for acyclic rules, bottom-up join orders (rooted arbitrarily) along a join tree, after early projections, usually yield tight intermediate size bounds.
A join spanning tree (JST) extends this notion to cyclic rules. A JST is a maximum spanning tree of the weighted join graph and reduces to a join tree when the rule is acyclic (Maier83). A rooted JST defines a join-project plan by post-order traversal: at each step, an atom is joined with its parent, followed by projecting out variables no longer needed. The only JST (also a join graph or a join tree) for Example 2.1 is shown in Figure 3. Here we root the JST at $\textsf{edge}(x,y)$, and it maps into the IR on the right, i.e. a bottom-up join order. If we root the JST at $\textsf{reach}(z)$ instead, we recover the IR in Figure 2(b).
@@
Search Space. Our search space excludes semijoins (atoms whose variables are subsumed by another), antijoins and filters, as they are pushed down to the lowest possible transformation in the IR. This narrows down to a multi-way (inner) join where the search space is defined as all rooted JSTs of its join graph. There are three reasons for such a choice: (1) JSTs avoid cross products when possible, since cross products correspond to zero-weight edges in the weighted join graph, (2) this space is reasonably small and easy to enumerate (Winter86), and (3) it collapses down to rooted join trees for acyclic rules, for which the post-order join-project plans yield tight bounds on the intermediate sizes (proven by (grounding24)). In our running example (which is acyclic), Figure 3 will be selected against Figure 2(b) as it has a cost of 2, instead of 3 in our cost model. Intuitively, it avoids computing $\textsf{edge}(x,y)\bowtie\textsf{edge}(y,z)$ by using two semi-joins.
@@
Consider an expensive rule (Figure 4, left) from the DOOP program analysis framework (BravenboerS09), which contains a recursive 8-way join. VarType, HeapAllocationType, and ComponentType are edbs; others (in bold) are idbs. Figure 4(center) shows its join graph. The JST selected by the optimizer is rooted at LoadArrayIdx and is annotated on the join graph by thick, directed edges. The dotted edges are edges in the join graph, but are not part of the rooted JST. Here, Reach(inm) is a semijoin atom to be subsumed by LoadArrayIdx and inm is projected away after semijoin. When creating the IR, this semijoin is pushed down to the leaf LoadArrayIdx.
@@ 5.3. Plan Execution
Left-deep Plans. Join planning is tied closely to the underlying execution. The listing order Soufflé, DDlog and others used is equivalent to a left-deep join plan in a dbms and is incapable of representing bushy plans, such as the one in Figure 4. This is a necessary choice for Soufflé as it always compiles the listing order into an indexed nested loop join, where indexes (e.g. B-trees) are built on every atom except the first one and it iterates over each tuple of the first atom (as the outer loop) and probes into the indexes of the rest. Left-deep plans allow pipelining without intermediate materialization, which makes Soufflé memory-efficient. However, scaling indexed nested loop joins to a multi-threaded setting is non-trivial (e.g. (LeisBK014)) and modern Datalog systems such as Soufflé and Flan (Flan24) only distribute the outermost for-loop among threads. As shown in our experiments, this level of parallelism is insufficient to saturate resources even for simple recursive queries like reachability.
Bushy Plans. FlowLog targets DD, a fundamentally different execution model. Its differential operators are inherently stateful, so intermediate materialization is unavoidable. A key advantage, however, is their asynchrony: changes propagate through the dataflow graph and operators react concurrently without waiting for earlier stages to finish. Hence, DD naturally exploits multicore parallelism, as computation is scheduled dynamically based on available updates rather than a pre-determined execution sequence. Consequently, FlowLog can take advantage of bushy plans (i.e. IR is bushy) without worrying about compromising parallelism. To control memory blow-ups, FlowLog leverages insights from database theory (e.g. tree decompositions (FAQ16; PANDA17; grounding24), worst-case optimal joins (NPRR13; freejoin24), etc.) to select plans that have smallest possible worst-case intermediate sizes. When multiple candidate plans have the same estimated cost, FlowLog heuristically prefers bushier (shallower) rooted JSTs.